How many electrons can be in the n = 6, l = 1 subshell?

Express your answer as an integer.
Answer #1

For a given L : the number of sub-shell = 2L + 1

So. For L = 1:

Number of sub-shell = (2)(1) + 1 = 3

In each shell maximum number of electrons = 2

Therefore: the total Number of electrons in L = 1 will be = (2)(3) = 6

Thus: Answer is 6